segment p1 p2 has length L and endpoints p1=(2,7)and p2=(8,3).
find a point on p1 p2 that is 1/3L away from p1.
A.(6,11/3)
B.(4,5)
C.(4,17/3)
D.(11/3,17/3)
E.n/a

Answer by stanbon(75887)

segment p1 p2 has length L and endpoints p1=(2,7)and p2=(8,3).
find a point on p1 p2 that is 1/3L away from p1.
------------
Let the point be (x,y):
EQUATION:
distance from (2,7) to (8,3) = sqrt(4^2+6^2) = sqrt(52)
---------
EQUATION:
distance from (2,7) to (x,y) = (1/3)(sqrt(52))
(y-7)^2 + (x-2)^2 = 52/9
---------------------------
The solution is a circle of points.
--------------------------
Cheers,
Stan H.